OpenStudy (anonymous):

can someone just check to see if this is correct? in the sequence that begins with: -6, 0, 6, 12... the 50th term is 288, correct? i did (49 x 6) + (-6) but I need to be 100% certain I am correct! thanks

OpenStudy (welshfella):

this is an arithmetic series with first term a = -6 and common difference d = 6 so 50th term = a + d(n - 1) = -6 + 6*49 = 288

OpenStudy (anonymous):

thanks! :)

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

6 x 48 = 288